A cannon launches a cannonball at an initial speed of 70.0 m/s, at an angle of 55° up from the horizontal. The ball strikes a castle wall 8.0 s later. How fast is the ball going when it strikes the wall?(ANSWER GIVEN:45 m/s)
Answer #1

vertical velocity after 8 s

vy = 70 sin 55 - 9.8* 8 = - 21.06 m/s

horizontal velocity

vx = 70 cos 55 = 40.15 m/s

net velocity at that instant

v^2 = vx^2 + Vy^2

v = 45.338 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck